### v1

# Should We Still Pretrain Encoders with Masked Language Modeling?

Abstract
Learning high-quality text representations is fundamental to a wide range of NLP tasks. While encoder pretraining has traditionally relied on Masked Language Modeling (MLM), recent evidence suggests that decoder models pretrained with Causal Language Modeling (CLM) can be effectively repurposed as encoders, often surpassing traditional encoders on text representation benchmarks. However, it remains unclear whether these gains reflect an inherent advantage of the CLM objective or arise from confounding factors such as model and data scale. In this paper, we address this question through a series of large-scale, carefully controlled pretraining ablations, training a total of 30 models ranging from 210 million to 1 billion parameters, and conducting over 15,000 fine-tuning and evaluation runs. We find that while training with MLM generally yields better performance across text representation tasks, CLM-trained models are more data-efficient and demonstrate improved fine-tuning stability. Building on these findings, we experimentally show that a biphasic training strategy that sequentially applies CLM and then MLM, achieves optimal performance under a fixed computational training budget. Moreover, we demonstrate that this strategy becomes more appealing when initializing from readily available pretrained CLM models (from the existing LLM ecosystem), reducing the computational burden needed to train best-in-class encoder models. We release all project artifacts at https://hf.co/MLMvsCLM to foster further research.
**footnotetext: Equal contribution
1 Introduction
Figure 1: Experimental setup overview and key results on sequence classification (610M model size, 40% MLM ratio). The upper plot shows downstream performance as a function of the CLM–to-MLM step ratio during pretraining (configurations outperforming MLM-only are highlighted in yellow). The lower plot illustrates the effect of applying MLM CPT to models initially trained with either MLM or CLM (Base).
Learning meaningful representations is essential for a wide range of NLP tasks, such as sequence classification, named entity recognition, extractive question answering, and information retrieval. Traditionally, these representations are learned with encoders pretrained solely with Masked Language Modeling (MLM) (Devlin et al., 2019). More recently, however, decoder models pretrained with Causal Language Modeling (CLM) and later adapted with MLM have challenged the MLM-only paradigm (BehnamGhader et al., 2024), achieving state-of-the-art results on the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2023; Enevoldsen et al., 2025). Yet, these results have so far only been observed on models that are significantly larger than typical encoders and trained on substantially more data (Lee et al., 2024; Meng et al., 2024; Kim et al., 2024; Muennighoff et al., 2024). As a result, it remains unclear whether the success of this new training paradigm stems from the causal objective itself, or merely from increased scale.
In this paper, we conduct a controlled study to investigate the impact of training with MLM and CLM objectives when learning textual representations. We compare models with equal architecture and size, trained on equal amount of data, and evaluate them on an extensive range of text representation tasks. We start by comparing pretraining exclusively on either objective, and then investigate a two-stage pretraining alternative, where training starts with the CLM objective and is followed by MLM. Additionally, given the availability of existing pretrained models, we investigate a continued pretraining scenario, where models are initialized from checkpoints trained exclusively with MLM or CLM, and then trained with MLM for a fraction of the pretraining steps. In total, we train 30 models ranging from 210 million to 1 billion parameters and conduct over 15,000 fine-tuning and evaluation runs, amounting to 110k MI250X GPU hours. Backed by these large-scale experiments, our results show that:
  • Although CLM training delivers strong performance on certain tasks and demonstrate good data efficiency and fine-tuning stability, MLM remains essential for achieving robust performance across all tasks, underscoring that bidirectional training remains indispensable (Section 3).
  • When pretraining a model from scratch, a two-stage protocol that first applies CLM followed by MLM offers a data-efficient way to build strong text representations by combining the strengths of both objectives (Section 4).
  • In a continued pretraining setting, adapting a CLM-pretrained model with MLM proves more effective than continuing MLM training from an MLM-pretrained model. This result suggests that leveraging widely pretrained decoder models is currently the best approach to obtain a strong encoder model (Section 5).
To facilitate reproducibility and support future research, we release all pretrained model checkpoints, [^1] https://huggingface.co/MLMvsCLM along with the training [^2] https://github.com/Nicolas-BZRD/EuroBERT/tree/MLM_vs_CLM and evaluation code. [^3] https://github.com/hgissbkh/EncodEval/tree/MLM_vs_CLM
2 Experimental Setup
In this section, we describe our controlled experimental setup, covering model architectures, pretraining and fine-tuning protocols, as summarized in Figure 1.
Models.
The model architectures closely follow those of the EuroBERT models (Boizard et al., 2025), with sizes of 210M, 610M, and 1B parameters. All models use a maximum context length of 2,048 tokens and a RoPE $\theta$ value of 10,000.444Additional details on the model architectures can be found in Appendix A.
Pretraining data.
Models are trained on unique English tokens from the FineWeb-Edu dataset (Penedo et al., 2024), which is known for supporting efficient model training. To ensure fair comparison across model configurations and training setups, all models are exposed to the same sequence of samples during training.
Pretraining objectives.
Models are trained using one of 3 approaches:
  1. 1.
    CLM uses next-token prediction, where each token is predicted autoregressively using a causal attention mask. The training objective is to minimize the negative log-likelihood:
[…]
    where $x=(x_{1},x_{2},\ldots,x_{T})$ is the input sequence and $P(x_{t}\mid x_{<t})$ is the model’s predicted distribution over the vocabulary given the preceding tokens.
  2. 2.
    MLM, by contrast, randomly masks a subset of tokens and trains the model to reconstruct them using a bidirectional attention mask. The objective is:
[…]
    where $\mathcal{M}\subset\{1,\ldots,T\}$ denotes the indices of masked tokens, and $x_{\setminus\mathcal{M}}$ is the input sequence with masked tokens replaced by special placeholders. We experiment with masking ratios of 20%, 30%, 40%, and 50%.
  3. 3.
    Finally, the two-stage CLM+MLM approach sequentially applies CLM pretraining followed by MLM.
Pretraining hyperparameters.
The default training setup uses 100 billion tokens, with a per-device batch size of 12 across 192 GPUs, resulting in an effective batch size of 2,359,296 tokens. We employ a Warmup-Stable-Decay (WSD) learning rate schedule: a 2,000-step warmup phase, followed by 38,000 steps with a constant learning rate of 5e-4, [^5] The optimal learning rate was selected from the range 1e-4 to 2e-3 based on experiments using 10% of the training data. ending with a 2,000-step linear decay phase, for a total of 42,000 training steps.
Pretraining setups.
To reflect two common real-world scenarios, we define two distinct pretraining setups, depending on whether the goal is to train a model from scratch or to continue training from an existing checkpoint:
  1. 1.
    Pretraining From Scratch (PFS): Models are trained from random initialization for a fixed number of steps using one of 3 objectives: CLM, MLM, or sequential CLM+MLM. For CLM and MLM, a standard WSD learning rate scheduler is applied. For CLM+MLM models, training is first performed with the CLM objective, and then resumed with MLM from CLM checkpoints that have not undergone learning rate decay.
  2. 2.
    Continued PreTraining (CPT): Models are initialized from existing checkpoints pretrained with either CLM or MLM, and training is resumed using the MLM objective. In contrast to the PFS setup, the pretrained models used for CPT have already undergone learning rate decay during their initial training phase, reflecting real-world constraints and continued training practices. Another key difference is that CPT starts from checkpoints where the loss has already converged, whereas in PFS, the objective switch typically occurs while gradient norms are still large and learning is active.
Fine-tuning tasks and datasets.
We evaluate all models across a broad range of text representation tasks, focusing on 4 key categories commonly used in real-world applications. For Sequence Classification (SC), we use SST-2 (Socher et al., 2013), MNLI (Williams et al., 2018), and QQP (Wang et al., 2017). For Token Classification (TC), we evaluate on the English subsets of CoNLL (Tjong Kim Sang & De Meulder, 2003), OntoNotes (Hovy et al., 2006), and UNER (Mayhew et al., 2024). Question Answering (QA) is assessed using SQuAD (Rajpurkar et al., 2016), SQuAD-v2 (Rajpurkar et al., 2018), and ReCoRD (Wang et al., 2019). For Information Retrieval (IR), we use MS MARCO (Bajaj et al., 2016), NQ (Kwiatkowski et al., 2019), and the English subset of MLDR (Chen et al., 2024) for long-context evaluation. [^6] Further details are provided in Appendix B.
Fine-tuning protocol.
To ensure a fair comparison, all models are fine-tuned using a consistent protocol. Each model is trained for up to 1,000 steps or one full epoch, whichever comes first, using a batch size of 32. To account for differences in model architecture and task requirements, we perform a grid search over 6 learning rates (1e-5, 2e-5, 5e-5, 1e-4, 2e-4, and 5e-4) for each model-dataset pair, with 10% warmup followed by linear decay. The learning rate yielding the best validation performance is selected. Fine-tuning employs bidirectional attention mask with task-specific loss functions: for SC, we use cross-entropy on mean-pooled token embeddings; TC and QA are trained using token-level cross-entropy; and for IR, we rely on the InfoNCE loss (Oord et al., 2018) with in-batch negatives, using mean pooling. The entire procedure is repeated across 5 random seeds to account for training variability. Fine-tuning is conducted on the in-domain training set, except for NQ and MLDR, for which training is performed on MS MARCO.
Evaluation protocol.
SC is assessed with accuracy, TC and QA with F1 score, and IR with NDCG@10. Results are averaged over the 5 random seeds, with 95% confidence intervals reported.
Infrastructure.
All pretraining and fine-tuning experiments are carried out on MI250X GPUs provided by the Adastra supercomputer, using the AMD-optimized EuroBERT codebase (Boizard et al., 2025).
Experimental scale.
Drawing reliable conclusions about model design choices typically requires scaling up experiments, as statistically sound results often require a large number of runs and sufficient training to reach minimal convergence (Hoffmann et al., 2022). We therefore carefully design our experiments to ensure robust and well-supported findings. Pretraining runs are executed across 24 nodes with 8 GPUs each (192 GPUs in total), while fine-tuning is performed on single GPUs. The total compute budget amounts to 110k GPU hours and is broken down as follows:
  • Base pretraining includes all 3 model sizes (210M, 610M, and 1B), trained on both CLM and MLM objectives. For MLM, 4 masking ratios (20%, 30%, 40%, and 50%) are explored, resulting in a total of 15 models, each trained for 42,000 steps (100B tokens), and accounting for 81k GPU hours.
  • PFS experiments explore various CLM-to-MLM step ratios (0%–100%, 25%–75%, 50%–50%, 75%–25%, and 100%–0%) across 3 training lengths (12,000, 22,000, and 42,000 steps), focusing on the 610M model size with a 40% masking ratio for MLM. To minimize redundant computation, we reuse checkpoints from the base pretraining runs whenever possible, applying learning rate decay only during the final 2,000 steps. In total, these runs result in 9 new models and account for 112,000 additional training steps and 16k GPU hours.
  • CPT experiments are performed on 610M models with a 40% masking ratio and varying training lengths (2,000, 12,000, and 22,000 steps) applied to both CLM- and MLM-pretrained models. To avoid redundant computation, training is resumed from shared checkpoints whenever possible. These runs yield 6 new models, totaling 26,000 training steps and 4k GPU hours.
  • Model evaluation involves fine-tuning a total of 30 final checkpoints: 5 models with 210M parameters, 5 with 1B, and 20 with 610M. In addition, we include 6 intermediate 610M checkpoints from CLM and MLM pretraining. Each model is fine-tuned on 12 datasets, across 6 learning rates and 5 random seeds, resulting in 15,120 fine-tuning runs and consuming 9k GPU hours.
3 Pretraining with CLM or MLM
Figure 2: MLM vs. CLM downstream performance, averaged across tasks and reported for all model sizes. For MLM, results correspond to a 40% masking ratio.
Figure 3: Task-wise downstream performance across different masking ratios for all model sizes.
Figure 4: Downstream performance as a function of pretraining steps for CLM and MLM objectives. Results are reported for 610M models, with a 40% masking ratio for MLM.
Figure 5: Impact of the fine-tuning learning rate on MLM- vs. CLM-pretrained models. Error bars indicate the standard deviation of metric scores across all seeds and learning rates between 1e-5 and 1e-4. Results are shown for 610M-parameter models, with a 40% masking ratio for MLM.
This section presents preliminary results comparing the MLM and CLM pretraining objectives in terms of downstream performance, data efficiency, and training and fine-tuning stability.
MLM generally outperforms CLM on text representation tasks.
As expected, bidirectional attention during pretraining tends to enhance representation quality on downstream tasks (Figure 2). The performance gap between models trained with MLM and those trained with CLM is especially noticeable on SC and QA tasks, and remains consistent across model sizes. The largest discrepancy is observed on QA, which appears particularly sensitive to the absence of bidirectional attention during pretraining. Interestingly, some task-specific trends emerge: for instance, the MLM–to-CLM gap widens with increasing model size on SC, but narrows on IR.
There is no universally optimal masking ratio.
The masking ratio is a key design choice when pretraining models with MLM. As shown in Figure 3, there is no single best ratio, as it depends on both model size and downstream task, making MLM pretraining a delicate balance. Larger models tend to benefit from higher masking ratios, consistent with prior findings from Wettig et al. (2023). Across tasks, IR datasets consistently prefer higher masking ratios regardless of model size. In contrast, for token-level tasks such as TC and QA, smaller models perform better with lower masking ratios. For larger models (610M and 1B), the performance curves exhibit a U-shape, indicating improved performance at both low and high masking ratios. [^7] Additional results in Appendix C further highlight the influence of masking ratio, showing substantial variation even among datasets within the same task category. To reduce computational overhead, most subsequent experiments report results on 610M models trained with a 40% masking ratio for MLM, which provides a strong overall compromise across tasks.
CLM models can perform competitively.
Interestingly, although CLM-pretrained models tend to underperform on SC, QA, and IR tasks, they achieve competitive results on TC, and even outperform MLM models at the 610M size (Figure 2). This suggests that causal pretraining can produce strong token-level representations, highlighting its potential despite the absence of bidirectional context.
CLM is more data-efficient than MLM in the early stages of training.
As shown in Figure 4, CLM outperforms MLM in downstream performance during the early phase of training (up to around step 5,000). However, as training progresses, MLM models tend to catch up and often surpass CLM, while CLM shows more limited gains in later steps. This suggests that although CLM saturates earlier, it enables more efficient representation learning in fewer steps. Notably, this makes CLM an appealing option for data-scarce scenarios, such as pretraining on low-resource languages, or simply as a warmup stage before MLM-based encoder training.
CLM-based pretraining improves fine-tuning stability.
A key challenge in fine-tuning models for text representation tasks is selecting optimal hyperparameters, particularly the learning rate, which can be sensitive to factors like model size, task type, or dataset scale, making exhaustive grid searches computationally expensive. As shown in Figure 5, models pretrained with CLM demonstrate lower sensitivity to learning rate choices than those pretrained with MLM. This indicates that CLM pretraining provides a more stable initialization for fine-tuning, facilitating more reliable performance and reducing the need for extensive hyperparameter tuning.
4 Two-Stage CLM+MLM Pretraining
Figure 6: Impact of two-stage CLM+MLM pretraining on downstream performance under different training budgets (12,000, 22,000, and 42,000 steps). The x-axis shows the percentage of CLM steps allocated in the first phase. Experiments are conducted on 610M models, with a 40% masking ratio during MLM training.
Figure 7: Comparison of downstream performance variability across different masking ratios (20%, 30%, 40%, 50%) for CLM and CLM+MLM pretraining configurations. Error bars indicate the standard deviation across fine-tuning seeds and masking ratios. Results are reported for 610M models, using 42,000 training steps for MLM and a schedule of 40,000 CLM steps followed by 2,000 MLM steps for CLM+MLM.
Building on Section 3, which shows that CLM outperforms MLM in TC performance, data efficiency, and training stability, we explore the benefits of applying CLM and MLM sequentially during pretraining.
Under fixed compute constraints, starting pretraining with CLM and continuing with MLM yields better results than MLM alone.
Motivated by the strong performance of CLM in learning text representations, we investigate the impact of two-stage pretraining objectives, starting with CLM and transitioning to MLM. We evaluate this under fixed compute budgets of 12,000, 22,000, and 42,000 training steps, using different splits: 100% MLM, 25%-75%, 50%-50%, 75%-25%, and 100% CLM. Interestingly, as shown in Figure 6, combining CLM and MLM consistently improves downstream performance compared to using MLM alone, though the effect varies by task and training budget. Overall, a split between 25%-75% and 50%-50% seems to provide the best balance.
CLM-based models exhibit lower sensitivity to masking ratio.
As shown in Figure 7, adapted CLM models show less variation in performance across different masking ratios compared to fully MLM-trained models. Initial CLM pretraining appears to stabilize model weights, making adaptation more robust to masking ratio choices and yielding more consistent downstream performance less sensitive to this design parameter.
5 Continued Pretraining from CLM and MLM Models
Figure 8: Impact of performing MLM CPT on either CLM- or MLM-pretrained models. CPT is conducted for 22,000 steps on 610M models with a 40% masking ratio, following 42,000 steps of initial pretraining.
Figure 9: MLM loss curves for CLM- and MLM-pretrained models across the 3 CPT compute budgets (2,000, 12,000, and 22,000 steps). Results are reported for 610M models, with MLM using a 40% masking ratio.
Figure 10: Downstream performance as a function of CPT length for CLM- and MLM-pretrained models, reported for settings of 2,000, 12,000, and 22,000 training steps. Experiments use 610M models with a 40% MLM ratio.
In this section, we focus on the CPT setting and analyze whether it is more effective to adapt a CLM-pretrained model using MLM or to continue MLM training from an MLM-pretrained model.
MLM CPT on a CLM-pretrained model outperforms MLM-only training.
We consider a setting in which equally-sized CLM and MLM models have been pretrained on the same data. The key question is whether it is more beneficial to spend additional compute on applying MLM CPT to the CLM model or to further train the MLM model. To keep computational costs manageable, we perform a 22,000-step CPT on the 610M model using a 40% masking ratio. As shown in Figure 8, the MLM-adapted CLM model consistently achieves superior downstream performance. On TC, where CLM-only models were already strong, performance is maintained and the gap to MLM remains. For QA and IR, the gap is effectively closed, while for SC, the MLM-adapted CLM model significantly outperforms the MLM-only model.
Fewer CPT steps already show strong performance.
Interestingly, we observe that it is not necessary to run as many as 22,000 CPT steps to achieve comparable performance between MLM and adapted CLM. As early as 12,000 steps, the results are already strong and broadly match those of MLM-only CPT in terms of loss (Figure 9) and downstream performance (Figure 10), with better results on TC and IR, comparable performance on SC, and nearly on par on QA. In terms of trends, applying MLM CPT on a CLM model also appears more promising, showing a steeper improvement curve toward the end, while MLM-only training seems to plateau (particularly noticeable on SC).
6 Related Work
Learning universal text representations has been a central focus in NLP, with MLM emerging as the dominant pretraining objective following the success of models like BERT (Devlin et al., 2019). MLM-based encoders leverage bidirectional attention to capture rich contextual dependencies, enabling strong performance across a wide range of tasks such as classification, question answering, and named entity recognition. Subsequent work has explored architectural extensions (e.g., RoBERTa (Liu et al., 2019), DeBERTa (He et al., 2021)) and more efficient pretraining strategies (Clark et al., 2020). In contrast, decoder-only architectures such as GPT-style models were originally optimized for autoregressive generation (Brown et al., 2020; Grattafiori et al., 2024; Jiang et al., 2023; Almazrouei et al., 2023), making them less naturally suited for text representation tasks.
Nevertheless, numerous studies have successfully explored adapting decoder-only models (typically larger and trained on more data than standard encoders) for learning high-quality text representations (Muennighoff, 2022; Lee et al., 2024; Meng et al., 2024; Zhang et al., 2025; Lee et al., 2025). As a result, these adapted decoders now rank among the top-performing models on prominent embedding benchmarks such as MTEB (Muennighoff et al., 2023; Enevoldsen et al., 2025). However, their strong performance does not disentangle the underlying factors (likely a combination of model scale and data regime) notably leaving open the question of whether causal pretraining itself plays a meaningful role in downstream representation quality.
Several recent studies have investigated the key factors influencing the performance of text representation models. BehnamGhader et al. (2024) propose a general framework for adapting decoder-only models to embedding tasks, accompanied by a series of ablations aimed at isolating the primary contributors to their performance. Wettig et al. (2023) explore the impact of the masking ratio on downstream tasks such as sequence classification and question answering, with a focus on learning dynamics, model scale, and fine-tuning behavior. Building on this line of work, we present a comprehensive analysis of how the choice of learning objective and attention mask during training influences downstream representation quality. By employing a unified and controlled pretraining setup, we ensure fair comparisons and minimize confounding factors, allowing for more precise attribution of the effects observed.
7 Conclusion and Future Work
Our study questions the established belief that Masked Language Modeling (MLM) is universally optimal for encoder pretraining. Based on large-scale experiments, we provide extensive empirical evidence indicating that encoders should not be pretrained exclusively with MLM objectives. Specifically, we show that decoder-only models trained with Causal Language Modeling (CLM) offer distinct advantages in terms of data efficiency, stability, and performance on certain text representation tasks. Importantly, our experiments demonstrate that adapting CLM models through subsequent MLM training consistently outperforms MLM training only, suggesting a nuanced interaction between both training strategies. We hope our findings and resource release will enable future work in this space.
Future work. Future research could explore extending our findings to Vision-Language Models (VLMs), many of which are decoder-based, potentially enhancing their representation learning capabilities. Additionally, future efforts could leverage our insights to pretrain state-of-the-art encoders, building on the combined strengths of CLM and MLM strategies identified in this study.
Limitations. Our study primarily focuses on large-scale ablations conducted exclusively in English. It remains uncertain whether our findings generalize to other languages, as linguistic and structural variations could influence the effectiveness of pretraining strategies. Additionally, our analysis does not account for the common practice of post-training decoder models on specialized datasets, such as those targeting reasoning or mathematical tasks. As such, we have not explored how this additional post-training phase might influence model behaviors or scaling laws relative to models trained from scratch using identical encoder data distributions. Although we observed notable trends up to the 1B parameter scale, future research should investigate whether these dynamics persist, diminish, or evolve at significantly larger model scales and training datasets, particularly considering post-training effects. Finally, in this work, we primarily focus on traditionally reported encoder evaluation settings. However, newer usage paradigms, such as late interaction (Khattab & Zaharia, 2020) or late chunking techniques (Günther et al., 2024; Conti et al., 2025) which heavily rely on bidirectional attention, may reveal even larger performance disparities.
Acknowledgments
We gratefully acknowledge the ADASTRA supercomputer (CINES) for providing technical support and high-performance computing (HPC) resources through projects C1615122, GDA2401, and CAD16490. We also thank the French government for its support through the France 2030 program, as part of the ArGiMi project. This work was further supported by the European Union’s Horizon Europe Research and Innovation Actions (UTTER, grant 101070631), the DECOLLAGE project (ERC-2022-CoG, grant 101088763), and the Portuguese Recovery and Resilience Plan via project C645008882-00000055 (Center for Responsible AI). Additional support was provided by FCT/MECI through national funds and, when applicable, co-funded EU funds under UID/50008: Instituto de Telecomunicações.
Appendix A Training Setup Details
This appendix provides additional information on the training parameters used in this study. Appendix A outlines both the architectural configurations and training settings.
Model Size 210M 610M 1B
Architecture
[…]
Embedding Dimension 768 1,152 1,728
FFN Dimension 3,072 4,096 5,120
Attention Heads 12 18 18
Key/Value Heads 12 6 6
Layer Normalization RMSNorm
RMSNorm $\epsilon$ $1\times 10^{-5}$
Activation Function SwiGLU
Maximum Context Length 2,048
Positional Embeddings RoPE
RoPE $\theta$ 10,000
Tokenizer LLaMA 3
Vocabulary Size 128,256
Training
[…]
Learning Rate $5\times 10^{-4}$
Scheduler WSD (linear decay)
Warmup Steps 2,000
Decay Steps 5%
Optimizer AdamW
Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.95
Adam $\epsilon$ $1\times 10^{-5}$
Weight Decay 0.1
Gradient Clipping Norm 1.0
Per-GPU Batch Size 12
Gradient Accumulation Steps 1
GPUs 192
Tokens/Step 2,359,296
Table 1: Architecture and training settings for the 3 model sizes under consideration.
Appendix B Details on Evaluation Datasets
This appendix offers additional details on the datasets used for evaluation.
SC datasets:
All SC datasets are sourced form the GLUE (Wang et al., 2018) benchmark.
  • MNLI (Williams et al., 2018) — MNLI is a large-scale benchmark for natural language inference, where each example pairs a premise with a hypothesis, and the task is to determine whether the hypothesis is entailed by, contradicts, or is neutral with respect to the premise. Since the test set is not publicly available, we report results on the validation set and reserve 5% of the training data for validation. We evaluate using the matched subset, which reflects in-domain performance.
  • QQP (Wang et al., 2017) — Binary classification dataset from Quora that asks whether a pair of questions are semantically equivalent. Similarly to MNLI, we report results on the validation set and reserve 5% of the training data for validation.
  • SST-2 (Socher et al., 2013) — Sentiment classification dataset consisting of movie review sentences. Each sentence is labeled as expressing either a positive or negative sentiment. We report results on the validation set and reserve 5% of the training data for validation.
TC datasets:
  • CoNLL (Tjong Kim Sang & De Meulder, 2003) — NER dataset comprising English and German newswire text, annotated with 4 entity types: person, organization, location, and miscellaneous. In this work, we use only the English portion.
  • OntoNotes (Hovy et al., 2006) — A large-scale dataset for NER and other NLP tasks, spanning diverse genres including newswire, broadcast news, and conversational speech in both English and Chinese. For this study, we focus on the English portion and evaluate on named entity spans.
  • UNER (Mayhew et al., 2024) — Multilingual NER dataset based on Wikipedia and Wikidata annotations. For this work, we use only the English subset.
QA datasets:
  • ReCoRD (Wang et al., 2019) — Challenging extractive QA dataset that tests commonsense reasoning by requiring models to select entities from a passage to answer cloze-style questions. Since the test set is not publicly available, we report results on the validation set and reserve 5% of the training data for validation.
  • SQuAD (Rajpurkar et al., 2016) — Benchmark for extractive question answering. It consists of questions posed on Wikipedia articles, with answers given as text spans within the corresponding passages. We report results on the validation set and reserve 5% of the training data for validation.
  • SQuAD-v2 (Rajpurkar et al., 2018) — Extension of SQuAD that includes both answerable and unanswerable questions. We report results on the validation set and reserve 5% of the training data for validation.
IR datasets:
  • MLDR (Chen et al., 2024) — Multilingual retrieval benchmark focused on long documents. In this work, we use only the English subset. Since MLDR does not provide training data, models are evaluated directly on the validation and test splits without task-specific fine-tuning.
  • MS MARCO (Bajaj et al., 2016) — Large-scale English passage retrieval dataset derived from real Bing search queries and web documents. MS MARCO is used in this work for both training and evaluation.
  • NQ (Kwiatkowski et al., 2019) — Open-domain English retrieval dataset based on real Google search queries, where answers are typically found in Wikipedia articles. As for MLDR, models are evaluated directly on the validation and test splits without task-specific fine-tuning.
To reduce computational overhead during evaluation on retrieval datasets, we restrict the corpus to labeled documents only, that is, those marked as positive or negative. Unlabeled documents are excluded from the evaluation.
Appendix C Detailed Results
C.1 MLM vs. CLM
This appendix provides detailed results comparing CLM and MLM objectives, including variations in masking ratios and model sizes. Section C.1, Section C.1, Section C.1, and Section C.1 present dataset-level results for SC, TC, QA, and IR, respectively.
Model Size Objective Masking SST-2 QQP MNLI Average
210M MLM 20% 89.93 ($\pm$0.26) 86.00 ($\pm$0.18) 79.29 ($\pm$0.32) 85.07 ($\pm$0.12)
30% 89.84 ($\pm$1.24) 85.82 ($\pm$0.36) 79.69 ($\pm$0.22) 85.12 ($\pm$0.40)
40% 90.44 ($\pm$0.71) 85.52 ($\pm$0.21) 78.72 ($\pm$0.20) 84.89 ($\pm$0.30)
50% 91.61 ($\pm$0.35) 85.52 ($\pm$0.49) 78.00 ($\pm$0.51) 85.04 ($\pm$0.25)
[…]
610M MLM 20% 92.61 ($\pm$0.57) 86.98 ($\pm$0.06) 83.08 ($\pm$0.33) 87.56 ($\pm$0.16)
30% 91.49 ($\pm$0.69) 86.48 ($\pm$0.25) 81.31 ($\pm$1.27) 86.43 ($\pm$0.51)
40% 91.42 ($\pm$0.62) 86.98 ($\pm$0.15) 82.59 ($\pm$0.16) 87.00 ($\pm$0.18)
50% 91.93 ($\pm$0.26) 86.38 ($\pm$0.34) 81.74 ($\pm$0.36) 86.68 ($\pm$0.13)
[…]
1B MLM 20% 92.39 ($\pm$0.47) 87.20 ($\pm$0.18) 84.33 ($\pm$0.13) 87.97 ($\pm$0.18)
30% 92.13 ($\pm$0.56) 86.46 ($\pm$0.20) 82.40 ($\pm$0.24) 87.00 ($\pm$0.13)
40% 92.89 ($\pm$0.22) 87.41 ($\pm$0.17) 84.40 ($\pm$0.25) 88.23 ($\pm$0.06)
50% 93.28 ($\pm$0.25) 87.27 ($\pm$0.17) 84.09 ($\pm$0.16) 88.21 ($\pm$0.12)
[…]
Table 2: MLM vs. CLM downstream performance results on SC datasets.
Model Size Objective Masking OntoNotes CoNLL UNER Average
210M MLM 20% 92.68 ($\pm$0.06) 91.38 ($\pm$0.15) 92.46 ($\pm$0.10) 92.17 ($\pm$0.09)
30% 92.71 ($\pm$0.14) 91.37 ($\pm$0.19) 92.45 ($\pm$0.13) 92.18 ($\pm$0.07)
40% 92.61 ($\pm$0.09) 91.70 ($\pm$0.19) 92.08 ($\pm$0.28) 92.13 ($\pm$0.11)
50% 92.46 ($\pm$0.12) 91.56 ($\pm$0.10) 92.25 ($\pm$0.16) 92.09 ($\pm$0.07)
[…]
610M MLM 20% 92.93 ($\pm$0.13) 92.34 ($\pm$0.17) 93.10 ($\pm$0.29) 92.79 ($\pm$0.09)
30% 92.69 ($\pm$0.15) 91.37 ($\pm$0.40) 91.94 ($\pm$0.64) 92.00 ($\pm$0.27)
40% 92.97 ($\pm$0.15) 91.58 ($\pm$0.44) 92.07 ($\pm$0.77) 92.21 ($\pm$0.26)
50% 92.95 ($\pm$0.20) 91.79 ($\pm$0.16) 92.84 ($\pm$0.33) 92.53 ($\pm$0.15)
[…]
1B MLM 20% 93.03 ($\pm$0.15) 92.28 ($\pm$0.30) 92.84 ($\pm$0.31) 92.72 ($\pm$0.19)
30% 92.93 ($\pm$0.17) 92.06 ($\pm$0.10) 92.55 ($\pm$0.16) 92.51 ($\pm$0.09)
40% 92.98 ($\pm$0.22) 92.05 ($\pm$0.14) 92.51 ($\pm$0.47) 92.51 ($\pm$0.16)
50% 93.12 ($\pm$0.25) 92.46 ($\pm$0.30) 93.06 ($\pm$0.55) 92.88 ($\pm$0.19)
[…]
Table 3: MLM vs. CLM downstream performance results on TC datasets.
Model Size Objective Masking SQuAD SQuAD-v2 ReCoRD Average
210M MLM 20% 74.66 ($\pm$0.26) 64.39 ($\pm$0.44) 16.15 ($\pm$6.01) 51.73 ($\pm$2.03)
30% 74.59 ($\pm$0.11) 61.87 ($\pm$1.02) 16.57 ($\pm$7.13) 51.01 ($\pm$2.58)
40% 73.73 ($\pm$0.11) 61.10 ($\pm$1.34) 13.45 ($\pm$2.93) 49.43 ($\pm$1.36)
50% 72.75 ($\pm$0.50) 58.47 ($\pm$1.96) 14.74 ($\pm$6.65) 48.65 ($\pm$2.99)
[…]
610M MLM 20% 76.97 ($\pm$0.19) 71.31 ($\pm$0.30) 46.74 ($\pm$1.11) 65.00 ($\pm$0.31)
30% 77.51 ($\pm$0.20) 65.98 ($\pm$2.56) 40.92 ($\pm$1.75) 61.47 ($\pm$0.61)
40% 77.76 ($\pm$0.29) 70.63 ($\pm$0.58) 39.93 ($\pm$5.26) 62.77 ($\pm$1.86)
50% 77.41 ($\pm$0.47) 68.61 ($\pm$0.53) 45.82 ($\pm$0.63) 63.95 ($\pm$0.47)
[…]
1B MLM 20% 77.81 ($\pm$0.57) 73.03 ($\pm$0.35) 54.38 ($\pm$0.92) 68.41 ($\pm$0.33)
30% 76.82 ($\pm$0.61) 66.09 ($\pm$2.13) 43.13 ($\pm$4.94) 62.02 ($\pm$1.84)
40% 79.93 ($\pm$0.27) 74.96 ($\pm$0.41) 55.95 ($\pm$0.99) 70.28 ($\pm$0.32)
50% 79.62 ($\pm$0.27) 73.53 ($\pm$0.46) 55.67 ($\pm$1.19) 69.61 ($\pm$0.56)
[…]
Table 4: MLM vs. CLM downstream performance results on QA datasets.
Model Size Objective Masking NQ MS MARCO MLDR Average
210M MLM 20% 80.23 ($\pm$0.54) 90.24 ($\pm$0.76) 53.76 ($\pm$0.90) 74.74 ($\pm$0.35)
30% 81.20 ($\pm$0.92) 90.44 ($\pm$0.75) 55.14 ($\pm$1.02) 75.59 ($\pm$0.57)
40% 81.43 ($\pm$1.35) 89.35 ($\pm$0.98) 56.56 ($\pm$0.62) 75.78 ($\pm$0.43)
50% 82.97 ($\pm$0.42) 89.31 ($\pm$0.35) 58.85 ($\pm$0.46) 77.04 ($\pm$0.27)
[…]
610M MLM 20% 85.14 ($\pm$0.49) 88.82 ($\pm$0.78) 57.07 ($\pm$0.93) 77.01 ($\pm$0.52)
30% 83.47 ($\pm$0.93) 90.36 ($\pm$0.81) 58.65 ($\pm$0.42) 77.50 ($\pm$0.42)
40% 86.76 ($\pm$0.63) 91.65 ($\pm$0.45) 60.25 ($\pm$0.76) 79.55 ($\pm$0.35)
50% 87.64 ($\pm$0.74) 90.73 ($\pm$1.26) 60.49 ($\pm$0.73) 79.62 ($\pm$0.54)
[…]
1B MLM 20% 86.02 ($\pm$0.55) 91.78 ($\pm$0.56) 59.85 ($\pm$0.40) 79.22 ($\pm$0.19)
30% 86.36 ($\pm$0.54) 90.90 ($\pm$0.61) 61.11 ($\pm$0.25) 79.46 ($\pm$0.04)
40% 87.89 ($\pm$0.66) 91.45 ($\pm$0.40) 60.51 ($\pm$0.80) 79.95 ($\pm$0.41)
50% 89.28 ($\pm$0.40) 92.98 ($\pm$0.65) 63.64 ($\pm$0.34) 81.97 ($\pm$0.27)
[…]
Table 5: MLM vs. CLM downstream performance results on IR datasets.
C.2 Data Efficiency
This appendix provides detailed results on data efficiency for MLM vs CLM models at different pretraining steps. Results are presented for 610M models with 40% masking for MLM. Table 6, Table 7, Table 8, and Table 9 present dataset-level results for SC, TC, QA, and IR, respectively.
Objective Training Step SST-2 QQP MNLI Average
MLM 1K 81.06 ($\pm$0.63) 73.78 ($\pm$0.18) 45.90 ($\pm$0.22) 66.91 ($\pm$0.26)
2K 81.40 ($\pm$0.81) 73.88 ($\pm$0.12) 49.40 ($\pm$0.33) 68.23 ($\pm$0.34)
5K 82.11 ($\pm$0.62) 78.44 ($\pm$1.06) 56.90 ($\pm$2.44) 72.49 ($\pm$1.17)
10K 86.90 ($\pm$1.93) 83.83 ($\pm$0.27) 73.37 ($\pm$1.62) 81.37 ($\pm$0.90)
20K 91.38 ($\pm$0.54) 85.97 ($\pm$0.28) 79.36 ($\pm$0.17) 85.57 ($\pm$0.20)
40K 90.94 ($\pm$0.67) 86.42 ($\pm$0.25) 81.84 ($\pm$0.30) 86.40 ($\pm$0.27)
CLM 1K 79.52 ($\pm$1.00) 73.56 ($\pm$0.24) 51.89 ($\pm$0.49) 68.32 ($\pm$0.28)
2K 84.24 ($\pm$0.47) 76.03 ($\pm$0.70) 60.54 ($\pm$0.89) 73.61 ($\pm$0.26)
5K 87.11 ($\pm$0.93) 81.80 ($\pm$0.31) 66.76 ($\pm$0.34) 78.56 ($\pm$0.46)
10K 89.33 ($\pm$0.29) 82.66 ($\pm$0.24) 69.75 ($\pm$0.22) 80.58 ($\pm$0.14)
20K 90.50 ($\pm$0.35) 83.74 ($\pm$0.08) 73.87 ($\pm$0.45) 82.70 ($\pm$0.07)
40K 90.94 ($\pm$0.43) 84.20 ($\pm$0.21) 75.31 ($\pm$0.43) 83.48 ($\pm$0.21)
Table 6: MLM vs. CLM downstream performance on SC datasets at various pretraining checkpoints. Results correspond to 610M models with MLM using 40% masking.
Objective Training Step OntoNotes CoNLL UNER Average
MLM 1K 84.39 ($\pm$0.45) 82.47 ($\pm$0.28) 82.58 ($\pm$0.78) 83.15 ($\pm$0.26)
2K 88.64 ($\pm$0.16) 86.02 ($\pm$0.19) 85.74 ($\pm$0.64) 86.80 ($\pm$0.23)
5K 90.93 ($\pm$0.07) 88.10 ($\pm$0.94) 89.05 ($\pm$0.89) 89.36 ($\pm$0.61)
10K 92.03 ($\pm$0.14) 90.60 ($\pm$0.22) 91.31 ($\pm$0.13) 91.31 ($\pm$0.08)
20K 92.62 ($\pm$0.07) 91.36 ($\pm$0.21) 92.06 ($\pm$0.12) 92.02 ($\pm$0.03)
40K 92.86 ($\pm$0.17) 91.45 ($\pm$0.28) 91.75 ($\pm$0.59) 92.02 ($\pm$0.19)
CLM 1K 88.09 ($\pm$0.13) 85.51 ($\pm$0.19) 87.44 ($\pm$0.35) 87.01 ($\pm$0.15)
2K 90.16 ($\pm$0.11) 88.82 ($\pm$0.21) 89.57 ($\pm$0.31) 89.52 ($\pm$0.15)
5K 91.54 ($\pm$0.10) 90.82 ($\pm$0.29) 91.15 ($\pm$0.28) 91.17 ($\pm$0.08)
10K 92.27 ($\pm$0.11) 91.47 ($\pm$0.18) 92.49 ($\pm$0.22) 92.08 ($\pm$0.10)
20K 92.69 ($\pm$0.08) 92.21 ($\pm$0.16) 92.99 ($\pm$0.30) 92.63 ($\pm$0.12)
40K 92.75 ($\pm$0.11) 92.25 ($\pm$0.14) 92.97 ($\pm$0.06) 92.65 ($\pm$0.03)
Table 7: MLM vs. CLM downstream performance on TC datasets at various pretraining checkpoints. Results correspond to 610M models with MLM using 40% masking.
Objective Training Step SQuAD SQuAD-v2 ReCoRD Average
MLM 1K 0.00 ($\pm$0.00) 50.07 ($\pm$0.00) 0.00 ($\pm$0.00) 16.69 ($\pm$0.00)
2K 0.94 ($\pm$0.50) 50.07 ($\pm$0.00) 0.00 ($\pm$0.00) 17.00 ($\pm$0.17)
5K 33.80 ($\pm$15.60) 49.87 ($\pm$0.27) 0.01 ($\pm$0.01) 27.89 ($\pm$5.17)
10K 59.19 ($\pm$12.12) 51.43 ($\pm$1.79) 4.38 ($\pm$2.42) 38.33 ($\pm$4.13)
20K 73.09 ($\pm$0.29) 58.38 ($\pm$0.63) 13.79 ($\pm$3.53) 48.42 ($\pm$1.27)
40K 77.00 ($\pm$0.27) 68.72 ($\pm$0.87) 27.98 ($\pm$10.00) 57.90 ($\pm$3.27)
CLM 1K 4.52 ($\pm$0.75) 50.00 ($\pm$0.09) 0.00 ($\pm$0.00) 18.17 ($\pm$0.24)
2K 30.93 ($\pm$11.46) 49.65 ($\pm$0.27) 0.10 ($\pm$0.05) 26.89 ($\pm$3.86)
5K 52.27 ($\pm$0.76) 50.52 ($\pm$0.49) 0.11 ($\pm$0.11) 34.30 ($\pm$0.11)
10K 62.31 ($\pm$0.51) 53.17 ($\pm$0.52) 0.32 ($\pm$0.07) 38.60 ($\pm$0.32)
20K 66.68 ($\pm$0.57) 54.99 ($\pm$0.62) 0.57 ($\pm$0.29) 40.75 ($\pm$0.19)
40K 69.40 ($\pm$0.38) 56.17 ($\pm$0.47) 0.41 ($\pm$0.08) 41.99 ($\pm$0.13)
Table 8: MLM vs. CLM downstream performance on QA datasets at various pretraining checkpoints. Results correspond to 610M models with MLM using 40% masking.
Objective Training Step NQ MS MARCO MLDR Average
MLM 1K 8.41 ($\pm$0.61) 43.97 ($\pm$2.88) 8.20 ($\pm$0.87) 20.19 ($\pm$1.38)
2K 21.64 ($\pm$0.82) 54.88 ($\pm$1.33) 15.15 ($\pm$0.97) 30.55 ($\pm$0.84)
5K 50.47 ($\pm$0.74) 78.14 ($\pm$0.59) 34.89 ($\pm$0.69) 54.50 ($\pm$0.44)
10K 69.32 ($\pm$0.90) 85.35 ($\pm$1.73) 44.30 ($\pm$0.57) 66.33 ($\pm$0.61)
20K 82.50 ($\pm$0.54) 89.60 ($\pm$0.65) 56.36 ($\pm$1.19) 76.16 ($\pm$0.62)
40K 85.46 ($\pm$0.24) 91.50 ($\pm$0.54) 60.83 ($\pm$0.29) 79.26 ($\pm$0.23)
CLM 1K 22.24 ($\pm$0.92) 54.20 ($\pm$1.70) 13.39 ($\pm$0.79) 29.94 ($\pm$0.49)
2K 45.12 ($\pm$0.68) 72.93 ($\pm$3.32) 25.26 ($\pm$0.87) 47.77 ($\pm$1.09)
5K 70.67 ($\pm$0.89) 85.19 ($\pm$0.81) 39.79 ($\pm$1.47) 65.22 ($\pm$0.60)
10K 78.47 ($\pm$1.94) 88.23 ($\pm$0.68) 46.14 ($\pm$1.36) 70.95 ($\pm$1.05)
20K 83.80 ($\pm$0.88) 89.49 ($\pm$0.75) 53.38 ($\pm$0.68) 75.56 ($\pm$0.45)
40K 84.88 ($\pm$0.71) 90.52 ($\pm$0.83) 52.76 ($\pm$1.65) 76.05 ($\pm$0.59)
Table 9: MLM vs. CLM downstream performance on IR datasets at various pretraining checkpoints. Results correspond to 610M models with MLM using 40% masking.
C.3 CLM-to-MLM Ratio
This appendix provides detailed results in the PFS setting, analyzing the impact of varying the CLM–MLM step ratio during pretraining. Results are presented for 610M models with 40% masking for MLM. Table 13, Table 11, Table 12, and Table 13 present dataset-level results for SC, TC, QA, and IR, respectively.
Total Steps CLM+MLM Mix SST-2 QQP MNLI Average
12K 0K+12K 89.33 ($\pm$0.68) 84.62 ($\pm$0.18) 75.17 ($\pm$0.33) 83.04 ($\pm$0.28)
3K+9K 90.09 ($\pm$0.45) 84.59 ($\pm$0.10) 76.32 ($\pm$0.36) 83.67 ($\pm$0.19)
6K+6K 90.02 ($\pm$0.34) 84.70 ($\pm$0.16) 76.15 ($\pm$0.32) 83.63 ($\pm$0.17)
9K+3K 89.77 ($\pm$0.77) 83.93 ($\pm$0.07) 73.25 ($\pm$0.15) 82.32 ($\pm$0.25)
12K+0K 89.66 ($\pm$0.51) 82.54 ($\pm$0.13) 70.40 ($\pm$0.38) 80.86 ($\pm$0.13)
22K 0K+22K 91.63 ($\pm$0.30) 86.30 ($\pm$0.08) 80.27 ($\pm$0.21) 86.07 ($\pm$0.07)
5K+17K 91.28 ($\pm$0.80) 86.42 ($\pm$0.32) 81.24 ($\pm$0.31) 86.31 ($\pm$0.29)
11K+11K 91.90 ($\pm$0.71) 86.69 ($\pm$0.12) 80.39 ($\pm$0.22) 86.33 ($\pm$0.25)
17K+5K 90.94 ($\pm$0.40) 85.86 ($\pm$0.16) 79.07 ($\pm$0.07) 85.29 ($\pm$0.18)
22K+0K 90.80 ($\pm$0.22) 83.86 ($\pm$0.17) 74.26 ($\pm$0.27) 82.97 ($\pm$0.15)
42K 0K+42K 91.42 ($\pm$0.62) 86.98 ($\pm$0.15) 82.59 ($\pm$0.16) 87.00 ($\pm$0.18)
10K+32K 92.41 ($\pm$0.18) 87.32 ($\pm$0.23) 83.83 ($\pm$0.25) 87.85 ($\pm$0.13)
21K+21K 92.36 ($\pm$0.39) 87.11 ($\pm$0.18) 83.28 ($\pm$0.14) 87.58 ($\pm$0.11)
32K+10K 92.16 ($\pm$0.43) 87.14 ($\pm$0.10) 81.90 ($\pm$0.24) 87.06 ($\pm$0.15)
42K+0K 91.10 ($\pm$0.39) 84.14 ($\pm$0.12) 75.49 ($\pm$0.22) 83.58 ($\pm$0.17)
Table 10: Impact of the CLM-to-MLM pretraining steps ratio on downstream performance in the PFS setup for SC datasets. Results are reported for 610M models with MLM using 40% masking.
Total Steps CLM+MLM Mix OntoNotes CoNLL UNER Average
12K 0K+12K 92.32 ($\pm$0.15) 90.81 ($\pm$0.22) 91.59 ($\pm$0.20) 91.58 ($\pm$0.10)
3K+9K 92.64 ($\pm$0.10) 91.63 ($\pm$0.09) 92.20 ($\pm$0.24) 92.16 ($\pm$0.11)
6K+6K 92.69 ($\pm$0.12) 91.62 ($\pm$0.20) 92.15 ($\pm$0.23) 92.15 ($\pm$0.09)
9K+3K 92.79 ($\pm$0.06) 91.83 ($\pm$0.26) 92.96 ($\pm$0.15) 92.53 ($\pm$0.10)
12K+0K 92.29 ($\pm$0.09) 91.67 ($\pm$0.25) 92.47 ($\pm$0.22) 92.14 ($\pm$0.06)
22K 0K+22K 92.83 ($\pm$0.15) 91.62 ($\pm$0.14) 92.39 ($\pm$0.24) 92.28 ($\pm$0.15)
5K+17K 93.09 ($\pm$0.08) 92.15 ($\pm$0.16) 93.08 ($\pm$0.53) 92.77 ($\pm$0.13)
11K+11K 93.15 ($\pm$0.04) 92.42 ($\pm$0.25) 92.93 ($\pm$0.30) 92.83 ($\pm$0.08)
17K+5K 93.13 ($\pm$0.12) 92.46 ($\pm$0.18) 92.86 ($\pm$0.24) 92.82 ($\pm$0.12)
22K+0K 92.73 ($\pm$0.06) 92.22 ($\pm$0.37) 92.94 ($\pm$0.52) 92.63 ($\pm$0.15)
42K 0K+42K 92.97 ($\pm$0.15) 91.58 ($\pm$0.44) 92.07 ($\pm$0.77) 92.21 ($\pm$0.26)
10K+32K 93.35 ($\pm$0.07) 92.39 ($\pm$0.20) 93.45 ($\pm$0.36) 93.07 ($\pm$0.15)
21K+21K 93.21 ($\pm$0.15) 92.43 ($\pm$0.13) 93.02 ($\pm$0.43) 92.89 ($\pm$0.15)
32K+10K 93.32 ($\pm$0.04) 92.41 ($\pm$0.17) 93.02 ($\pm$0.24) 92.92 ($\pm$0.11)
42K+0K 92.79 ($\pm$0.05) 92.16 ($\pm$0.21) 93.12 ($\pm$0.29) 92.69 ($\pm$0.14)
Table 11: Impact of the CLM-to-MLM pretraining steps ratio on downstream performance in the PFS setup for TC datasets. Results are reported for 610M models with MLM using 40% masking.
Total Steps CLM+MLM Mix SQuAD SQuAD-v2 ReCoRD Average
12K 0K+12K 67.95 ($\pm$0.65) 53.73 ($\pm$2.04) 2.85 ($\pm$1.06) 41.51 ($\pm$0.55)
3K+9K 71.36 ($\pm$0.63) 56.28 ($\pm$0.44) 3.68 ($\pm$1.81) 43.77 ($\pm$0.63)
6K+6K 71.29 ($\pm$0.27) 55.85 ($\pm$0.89) 1.19 ($\pm$0.40) 42.78 ($\pm$0.37)
9K+3K 68.20 ($\pm$0.83) 53.85 ($\pm$1.09) 2.55 ($\pm$4.07) 41.53 ($\pm$1.81)
12K+0K 63.40 ($\pm$0.72) 53.49 ($\pm$0.40) 0.32 ($\pm$0.11) 39.07 ($\pm$0.34)
22K 0K+22K 74.02 ($\pm$0.38) 61.02 ($\pm$0.92) 19.29 ($\pm$5.45) 51.44 ($\pm$1.81)
5K+17K 75.71 ($\pm$0.93) 61.50 ($\pm$2.71) 26.08 ($\pm$6.25) 54.43 ($\pm$1.51)
11K+11K 75.43 ($\pm$0.21) 61.33 ($\pm$0.73) 31.14 ($\pm$6.92) 55.97 ($\pm$2.43)
17K+5K 74.15 ($\pm$0.46) 59.03 ($\pm$1.53) 26.99 ($\pm$3.43) 53.39 ($\pm$1.31)
22K+0K 67.52 ($\pm$0.40) 55.44 ($\pm$0.66) 0.47 ($\pm$0.14) 41.14 ($\pm$0.25)
42K 0K+42K 77.76 ($\pm$0.29) 70.63 ($\pm$0.58) 39.93 ($\pm$5.26) 62.77 ($\pm$1.86)
10K+32K 79.20 ($\pm$0.22) 72.33 ($\pm$0.81) 51.79 ($\pm$1.50) 67.77 ($\pm$0.53)
21K+21K 77.12 ($\pm$0.28) 67.81 ($\pm$0.66) 33.60 ($\pm$11.35) 59.51 ($\pm$3.84)
32K+10K 76.54 ($\pm$0.31) 64.49 ($\pm$0.82) 43.20 ($\pm$1.29) 61.41 ($\pm$0.66)
42K+0K 69.42 ($\pm$0.30) 56.44 ($\pm$0.10) 0.42 ($\pm$0.39) 42.09 ($\pm$0.20)
Table 12: Impact of the CLM-to-MLM pretraining steps ratio on downstream performance in the PFS setup for QA datasets. Results are reported for 610M models with MLM using 40% masking.
Total Steps CLM+MLM Mix NQ MS MARCO MLDR Average
12K 0K+12K 74.15 ($\pm$1.46) 85.95 ($\pm$1.83) 49.59 ($\pm$0.56) 69.90 ($\pm$1.21)
3K+9K 80.08 ($\pm$0.44) 89.63 ($\pm$0.87) 55.09 ($\pm$0.56) 74.94 ($\pm$0.22)
6K+6K 80.67 ($\pm$0.61) 88.30 ($\pm$0.79) 56.70 ($\pm$0.58) 75.22 ($\pm$0.41)
9K+3K 80.10 ($\pm$0.44) 89.23 ($\pm$1.28) 55.63 ($\pm$0.87) 74.99 ($\pm$0.31)
12K+0K 80.16 ($\pm$1.16) 88.88 ($\pm$1.27) 49.26 ($\pm$0.54) 72.77 ($\pm$0.82)
22K 0K+22K 83.72 ($\pm$0.85) 90.42 ($\pm$0.38) 58.47 ($\pm$0.45) 77.54 ($\pm$0.28)
5K+17K 84.07 ($\pm$0.52) 89.38 ($\pm$0.77) 58.42 ($\pm$1.00) 77.29 ($\pm$0.46)
11K+11K 84.48 ($\pm$0.50) 89.72 ($\pm$0.56) 58.22 ($\pm$0.47) 77.47 ($\pm$0.40)
17K+5K 83.36 ($\pm$0.57) 90.97 ($\pm$0.56) 59.12 ($\pm$0.26) 77.81 ($\pm$0.18)
22K+0K 84.01 ($\pm$0.75) 90.36 ($\pm$0.98) 52.75 ($\pm$2.03) 75.71 ($\pm$0.34)
42K 0K+42K 86.76 ($\pm$0.63) 91.65 ($\pm$0.45) 60.25 ($\pm$0.76) 79.55 ($\pm$0.35)
10K+32K 86.79 ($\pm$0.28) 91.98 ($\pm$1.35) 60.86 ($\pm$1.77) 79.88 ($\pm$0.87)
21K+21K 86.80 ($\pm$0.86) 90.98 ($\pm$0.92) 61.22 ($\pm$1.27) 79.67 ($\pm$0.81)
32K+10K 86.82 ($\pm$0.49) 91.40 ($\pm$0.92) 61.14 ($\pm$0.84) 79.79 ($\pm$0.36)
42K+0K 85.57 ($\pm$0.34) 89.88 ($\pm$1.17) 53.16 ($\pm$0.42) 76.20 ($\pm$0.54)
Table 13: Impact of the CLM-to-MLM pretraining steps ratio on downstream performance in the PFS setup for IR datasets. Results are reported for 610M models with MLM using 40% masking.
C.4 Continued Pretraining
This appendix presents detailed results for the CPT setting, where different CPT lengths are evaluated. All results correspond to 610M-parameter models trained with a 40% masking ratio under the MLM objective. Table 14, Table 15, Table 16, and Table 17 report dataset-level performance for SC, TC, QA, and IR, respectively.
PT Objective CPT Steps SST-2 QQP MNLI Average
MLM - 91.42 ($\pm$0.62) 86.98 ($\pm$0.15) 82.59 ($\pm$0.16) 87.00 ($\pm$0.18)
2K 91.54 ($\pm$0.54) 86.84 ($\pm$0.16) 82.85 ($\pm$0.27) 87.08 ($\pm$0.20)
12K 92.71 ($\pm$0.37) 86.89 ($\pm$0.31) 83.23 ($\pm$0.29) 87.61 ($\pm$0.21)
22K 92.16 ($\pm$0.59) 86.94 ($\pm$0.10) 83.31 ($\pm$0.35) 87.47 ($\pm$0.24)
CLM - 91.10 ($\pm$0.39) 84.14 ($\pm$0.12) 75.49 ($\pm$0.22) 83.58 ($\pm$0.17)
2K 92.06 ($\pm$0.48) 85.93 ($\pm$0.15) 78.18 ($\pm$0.23) 85.39 ($\pm$0.17)
12K 92.98 ($\pm$0.34) 87.48 ($\pm$0.13) 83.05 ($\pm$0.18) 87.84 ($\pm$0.15)
22K 93.62 ($\pm$0.33) 87.56 ($\pm$0.08) 84.02 ($\pm$0.10) 88.40 ($\pm$0.11)
Table 14: Impact of continued MLM pretraining on both MLM- and CLM-pretrained models across different sequence lengths on SC datasets. Results are reported for 610M models with 40% masking for MLM.
PT Objective CPT Steps OntoNotes CoNLL UNER Average
MLM - 92.97 ($\pm$0.15) 91.58 ($\pm$0.44) 92.07 ($\pm$0.77) 92.21 ($\pm$0.26)
2K 92.85 ($\pm$0.26) 90.89 ($\pm$1.10) 92.74 ($\pm$0.39) 92.16 ($\pm$0.41)
12K 93.02 ($\pm$0.17) 91.19 ($\pm$0.82) 92.59 ($\pm$0.51) 92.27 ($\pm$0.41)
22K 92.82 ($\pm$0.45) 91.41 ($\pm$0.16) 92.10 ($\pm$0.57) 92.11 ($\pm$0.32)
CLM - 92.79 ($\pm$0.05) 92.16 ($\pm$0.21) 93.12 ($\pm$0.29) 92.69 ($\pm$0.14)
2K 92.91 ($\pm$0.07) 92.71 ($\pm$0.26) 93.02 ($\pm$0.26) 92.88 ($\pm$0.13)
12K 92.78 ($\pm$0.56) 92.53 ($\pm$0.13) 93.01 ($\pm$0.29) 92.77 ($\pm$0.12)
22K 92.85 ($\pm$0.13) 92.30 ($\pm$0.19) 92.94 ($\pm$0.11) 92.70 ($\pm$0.05)
Table 15: Impact of continued MLM pretraining on both MLM- and CLM-pretrained models across different sequence lengths on TC datasets. Results are reported for 610M models with 40% masking for MLM.
PT Objective CPT Steps SQuAD SQuAD-v2 ReCoRD Average
MLM - 77.76 ($\pm$0.29) 70.63 ($\pm$0.58) 39.93 ($\pm$5.26) 62.77 ($\pm$1.86)
2K 77.92 ($\pm$0.21) 71.07 ($\pm$0.85) 41.69 ($\pm$4.59) 63.56 ($\pm$1.77)
12K 78.71 ($\pm$0.23) 72.35 ($\pm$0.67) 48.66 ($\pm$2.17) 66.57 ($\pm$0.85)
22K 79.01 ($\pm$0.18) 72.77 ($\pm$0.47) 51.21 ($\pm$1.32) 67.66 ($\pm$0.53)
CLM - 69.42 ($\pm$0.30) 56.44 ($\pm$0.10) 0.42 ($\pm$0.39) 42.09 ($\pm$0.20)
2K 73.48 ($\pm$0.28) 59.13 ($\pm$0.21) 10.94 ($\pm$5.90) 47.85 ($\pm$2.04)
12K 77.25 ($\pm$0.29) 65.62 ($\pm$1.15) 46.75 ($\pm$1.38) 63.21 ($\pm$0.17)
22K 78.13 ($\pm$0.27) 69.69 ($\pm$0.56) 52.06 ($\pm$0.72) 66.62 ($\pm$0.34)
Table 16: Impact of continued MLM pretraining on both MLM- and CLM-pretrained models across different sequence lengths on QA datasets. Results are reported for 610M models with 40% masking for MLM.
PT Objective CPT Steps NQ MS MARCO MLDR Average
MLM - 86.76 ($\pm$0.63) 91.65 ($\pm$0.45) 60.25 ($\pm$0.76) 79.55 ($\pm$0.35)
2K 86.77 ($\pm$0.84) 92.18 ($\pm$0.63) 62.31 ($\pm$0.42) 80.42 ($\pm$0.27)
12K 87.01 ($\pm$0.52) 92.32 ($\pm$0.92) 60.44 ($\pm$1.02) 79.92 ($\pm$0.47)
22K 87.58 ($\pm$0.62) 92.52 ($\pm$0.79) 61.26 ($\pm$0.53) 80.45 ($\pm$0.27)
CLM - 85.57 ($\pm$0.34) 89.88 ($\pm$1.17) 53.16 ($\pm$0.42) 76.20 ($\pm$0.54)
2K 86.46 ($\pm$0.38) 88.80 ($\pm$0.76) 60.45 ($\pm$0.43) 78.57 ($\pm$0.22)
12K 88.29 ($\pm$0.26) 89.56 ($\pm$0.92) 62.97 ($\pm$0.52) 80.27 ($\pm$0.29)
22K 87.67 ($\pm$0.93) 90.71 ($\pm$0.23) 63.72 ($\pm$0.41) 80.70 ($\pm$0.26)
Table 17: Impact of continued MLM pretraining on both MLM- and CLM-pretrained models across different sequence lengths on IR datasets. Results are reported for 610M models with 40% masking for MLM.